OpenStudy (anonymous):

Find average rate of change of the function over the given intervals. h(t)= cot t {5π/4, 7π/4}

OpenStudy (tkhunny):

Find the values of the given function at the endpoints of the interval.

zepdrix (zepdrix):

Yah do that ^ \[\Large\rm h\left(\frac{7\pi}{4}\right)=\cot\frac{7\pi}{4}\]\[\Large\rm h\left(\frac{7\pi}{4}\right)=?\]

OpenStudy (anonymous):



OpenStudy (tkhunny):

You could find the sine and cosine at the same place.

OpenStudy (anonymous):

Oh yeah! Thank you!

OpenStudy (anonymous):

For my answer I got -4/pi